OpenStudy (anonymous):

i need help in finding the solution of this ordinary differential equation 3(3x^2+y^2)dx-2xydy=0 using any method.

OpenStudy (goformit100):

@yeyenunez "Welcome to OpenStudy. I can answer your questions or guide you. Please use the chat for off topic questions. And remember to give the helper a medal, by clicking on "Best Answer". We follow a code of conduct, ( http://openstudy.com/code-of-conduct ). Please take a moment to read it."

OpenStudy (anonymous):

Lets answer :)

OpenStudy (anonymous):

@E.ali ok :D

OpenStudy (mimi_x3):

\[ 3(3x^2+ y^2)dx - 2xy\frac{dy}{dx}=0\] \[ 9x^2 + 3y^2dx - 2xy\frac{dy}{dx}=0\] \[ -2xy\frac{dy}{dx} + 3y^2dx + 9x^2 =0 => -2xy\frac{dy}{dx} + 3y^2dx = -9x^2\] \[ 2ydy - \frac{3y^2}{x} = 9x\] Let \(u = y^2\) \[ \frac{du}{dx} = 2y\frac{dy}{dx}\] \[ => \frac{du}{dx} - \frac{3u}{x} =9x\] so it's a linear ode where the solution is \[ y = e^{-h(x)} [ \int e^{h(x)rdx +c\]

OpenStudy (anonymous):

hint, equation is exact

OpenStudy (mimi_x3):

i dont think so..

OpenStudy (anonymous):

ups, you right. Didn't see the - sign

OpenStudy (anonymous):

\[3(3x ^{2}+y ^{2})dx-2xydy=0\] \[9x ^{2}+3y ^{2} -2xydy=0\] \[9dx-\frac{ 2xydy-3y ^{2}dx }{ x ^{2} }=0\] There is an integrable combination wherein: \[d(\frac{ y ^{2} }{ x })=\frac{ 2xydy-y ^{2}dx }{ x ^{2} }\] However I am having problems with the \[-3y ^{2}\] in order to apply that integrable combination :)

OpenStudy (anonymous):

this is homogeneous ode. Write it like this: \(y'=\huge\frac{2xy}{3x^2+y^2}\) now reduce it single variable by multiplaying up and down part by \(1/x^2\) . Later you can solve it by taking variable change \(u=y/x\)

OpenStudy (anonymous):

which makes it separate variable, and simple integration, :)

OpenStudy (anonymous):

@Mimi_x3 u'r solution is wrong

OpenStudy (mimi_x3):

how lol?

OpenStudy (anonymous):

where that 3 come out from?

OpenStudy (mimi_x3):

the question lol?

OpenStudy (anonymous):

@myko uhm its \[3(3x ^{2}+y ^{2})\] so if you get its partial derivatives its not homogeneous and there is also a negative sign O.O

OpenStudy (anonymous):

Hey guys ! I have a better way !:)

OpenStudy (mimi_x3):

is the ode exact?

OpenStudy (anonymous):

equation is homogenious

OpenStudy (anonymous):

and u solve the way i explaned

OpenStudy (anonymous):

@: myko : Wait please !

terenzreignz (terenzreignz):

Bernoulli?

OpenStudy (anonymous):

We have : 9x^2+2y^2.dx -2(x-y^2-d)=0 ok ?!

OpenStudy (anonymous):

\[f(x,y) = 3(3x ^{2}+y ^{2})dx - 2xydy=0\] \[f(\lambda x,\lambda y) = 3(3(\lambda x)^{2}+(\lambda y)^{2})- 2(\lambda x)(\lambda y) =0\] so it is homogenous lol =))))

OpenStudy (anonymous):

And now we can have : 9x^2+2y^2.dx=2(-x+y^2+d) Ok?!

OpenStudy (anonymous):

And now :! CAN you complate it ?!

OpenStudy (anonymous):

@E.ali can you use the equation button I can't understand your equation...

OpenStudy (anonymous):

the idea is to first separate the variables all y's on one side all x's on the other side to accomplish that, let's start by distributing in the first parenthesis go ahead and give that a try please.

OpenStudy (anonymous):

@ ; Archie : We dont have just x and y .! We have d to !

OpenStudy (anonymous):

distributing means multiplying the factor in front of the parentheses by each and every term inside the parenthesis here is an example: a(x+y) becomes ax + ay that is called distributing the a to the two terms inside the parenthesis.

OpenStudy (anonymous):

Sure we do, but we are trying to do this step by step. if you want, you can distribute both the 3 and the dx at the same time

OpenStudy (anonymous):

but I suggest you do it step by step

OpenStudy (anonymous):

@ⒶArchie☁✪ \[9x ^{2}dx+3y ^{2}dx=2xy dy\] \[9x ^{2}dx=2xy dy-3y ^{2}dx\]

OpenStudy (anonymous):

am i doing it right?

OpenStudy (anonymous):

Perfect :)

OpenStudy (anonymous):

but you don't need to move the 3y^2dx term over to the right let's keep that on the left, together with the other term that has dx in it have a look:

OpenStudy (anonymous):